# The Nuclear Star Cluster of M 74: a fossil record of the very early stages of a star-forming galaxy

### v1

@@ 5.1 Stellar populations of the galactic center: the NSC compared with the host galaxy
On the other hand, the difference between the mass-weighted ages of the NSC and its host galaxy is smaller (than when considering light-weighted ages), with the stars in the host galaxy being about 2 Gyr younger than the NSC. This is a significant difference for mass-weighted results, which are expected to be biased towards older ages (e.g., SanchezBlazquez2014a; Sattler2023; Sattler2025). Mass-weighted and light-weighted metallicities are quite similar to each other, while light-weighted [Mg/Fe] abundances are slightly higher than the mass-weighted ones. Uncertainties indicated in Table 1 were obtained via Monte Carlo simulations (Sect. 4.2). They represent only statistical uncertainties and they are relatively low for such high S/N spectra. Differences in the stellar-population properties of the different components are larger than these uncertainties. Systematic uncertainties, such as those due to imperfections in the SSP models, are very challenging to estimate and are not taken into account here.
As a reference, we can compare the results above with the stellar populations from the spectrum extracted from the original integrated data cube (in a PSF size, blue solid line in Fig. 4). This contains the integrated light of all structures in the LOS (the NSC, the disk and the bulge, see Sect. 4.1) and, as expected, gives intermediate stellar populations between the NSC and the host galaxy (Table 1). Finally, we can also compare the stellar populations of the NSC and of the galaxy in a central PSF aperture with the surrounding region of the galaxy. For that, we calculated the average ages, [M/H] and [Mg/Fe] within a much larger central aperture of 100 pc (excluding the central $3\sigma$ PSF to make sure the contamination from the NSC is negligible). For this, we used the stellar-population maps extracted from the original integrated data cube (Appendix A). Stars are younger (5.6 Gyr) and slightly more metal-rich (0.1-0.1 dex), than the stars in the central PSF aperture extracted from the same original data cube (Table 1). However, the [Mg/Fe] abundance is very similar (0.1 dex). The fact that the original data cube, including the light of stars of both the NSC and the host galaxy, shows gradients with older and more metal-poor stars in the very center, is in perfect agreement with the presence of an older and more metal-poor component in the center (i.e. the NSC), which is not present in the surroundings.
We have also tested the robustness of this gradient by fitting a spectrum from the original data cube, but integrated now in a $3\sigma$ PSF (about 40 pc, with a much stronger contribution from the host galaxy). In a $3\sigma$ PSF aperture, we find ages about 1 Gyr younger in the original and host-galaxy data cubes, while the NSC age remains the same. Since we are comparing results from slightly different methods (on the one hand, fitting spectra integrated in specific apertures, on the other hand, averaging stellar-population properties in the spatially resolved maps shown in Appendix A), we have checked that the same results were found within $1\sigma$ PSF with the different methods. The mean age, [M/H] and [Mg/Fe] from a $1\sigma$ PSF aperture in the spatially resolved maps in Fig. 8 are respectively: $6.8$ Gyr, 0.2-0.2 and $0.1$ dex, approximately the same as the results from fitting a spectrum from the original data cube obtained integrating in a $1\sigma$ PSF aperture (shown in Table 1).
5.2 Spatially resolved stellar populations of the host galaxy
We mapped the mean age, [M/H] and [Mg/Fe] of the host galaxy, fitting the host-galaxy data cube decomposed with C2D (Sect. 4.1.2 and 4.2). The maps are shown in Fig. 5. Both [M/H] and [Mg/Fe] are presented in logarithmic scale and with respect to solar values (zero values correspond to solar values). The age map (top panel) shows a relatively young host galaxy (with most regions younger than ${\sim}4$ Gyr). A pattern following the spiral arms can be identified, with alternative regions dominated either by young (about ${\sim}2$ Gyr-old or younger) or old stars (older than ${\sim}8$ Gyr). The [M/H] map (middle panel) displays a clear negative radial gradient, with a central metal-rich region of about 500 pc size (supersolar metallicities). Outer regions have alternatively higher and lower values. [Mg/Fe] abundances are mostly anti-correlated with [M/H], showing the lowest values (close to solar) in the central metal-rich ${\sim}500$ pc. Hints of the spiral-arm shape are visible in both [M/H] and [Mg/Fe] maps, whose values do not necessarily follow any correlation or anti-correlation with age. However, the youngest (outer) regions in the analyzed data cube show a strong $\alpha$ enhancement in some metal-poor blobs.
[…]
@@ 5.2 Spatially resolved stellar populations of the host galaxy
In Appendix A, we present for reference the analysis of the original MUSE data cube before the spectro-photometric decomposition. Although the Voronoi binning is slightly different, Fig. 8 shows similar trends to Fig. 5. The difference between the original data cube and the host-galaxy decomposed data cube is that the former includes the light of the NSC, which was extracted from the latter. Thus, we can expect a difference in the stellar populations only in the region where the NSC dominates. Indeed, as described in Appendix A, a clear [M/H] drop is visible within a PSF size in Fig. 8. Our results for the central region of M 74 are in agreement with the stellar-population analysis by Pessa2023 (Pessa2023, see also Sect. 3).
5.3 Star formation histories (SFHs) of the NSC and comparison with the host galaxy
We extracted SFHs of the NSC and the host galaxy from spectra extracted in our $1\sigma$ PSF aperture, in terms of mass fraction assigned to SSP models of different ages during the fitting process. We used the mass-weighted results for this purpose. We show these SFHs in Fig. 6, respectively in the top and middle panels. In the bottom panel we show, as a reference, the SFH of the full galaxy, extracted from the same PSF aperture in the original MUSE data cube before the C2D decomposition (blue spectrum in Fig 4).
The NSC shows different stellar populations at old ages, with the highest mass fractions distributed in ages between 10 and 11 Gyr. The host galaxy shows an old component made of two peaks of different ages (${\sim}10$ and ${\sim}12$ Gyr) and a relatively recent component with ages of about 2 Gyr and younger. Some intermediate-age populations are also present (${\sim}6$ Gyr). The SFH of the full galaxy shows a combination of all these components. A word of caution is deserved by the large uncertainties, indicated by the shades in each panel of Fig. 6. These shades cover the region between the 16% and 84% percentiles, calculated on the distribution of the 100 Monte Carlo realizations. Since mass fractions are normalized to the total mass in each SFH (in each fitted spectrum), these errors indicate that some mass fraction assigned to the highest peaks could be, in reality, redistributed to minor peaks. Overall, the SFHs suggest that the center of the galaxy had an intense early phase of star formation, between 14 and 8 Gyr ago, when the NSC stars formed. Afterwards, it experienced a long quiescent phase, with a very low star-formation rate. Only relatively recently (later than 3 Gyr ago), a more active star formation was reactivated in M 74. However, while this episode involved an extended region of the disk and was enhanced in spiral arms, as shown in Fig. 5, it did not affect the NSC, as this remained old and metal poor. SanchezBlazquez2014b found a very similar SFH for M 74, using the STEllar Content via Maximum A Posteriori likelihood (STECKMAP) spectral-fitting code (Ocvirk2006).
@@ 6.1 The nuclear star cluster in M 74 and its formation
Our stellar population analysis has shown an ancient and metal-poor NSC, surrounded by a much younger and more metal-rich region of the host galaxy. As shown by JWST observations (Hoyer2023b), the NSC lives in a cavity with no gas or dust, while being hosted by an actively star-forming galaxy with a complex structure of gas and dust filaments. The cavity formed due to some processes either keeping gas and dust out of it or wiping them out of the central region (see also Sect. 6.3). While the presence of a massive black hole in M 74 is still under debate, the presence of a low-luminosity AGN was proposed by different authors (e.g., Dong2006; She2017). It is also one of the possible explanations for the presence of the cavity (all gas may have been blown via AGN feedback, especially if the black-hole activity was much stronger in the past) and of the mid-infrared emission found at the center of the galaxy with JWST observations (Hoyer2023b). Similarly, central quiescent regions (of about 100 pc sizes), surrounded by starbursts, as well as nuclear older ages (within radii from few to ${\sim}50$ pc) than the circumnuclear regions, were previously detected in AGN hosts (e.g., Marquez2025; daSilva2025).
In any case, the presence of the cavity may have affected the evolution of the NSC, which could not grow via in-situ star formation due to the lack of gas. However, the surrounding region of the NSC, including stars in the region of the cavity, is much younger and more metal-rich. Our decomposition, supported by the analysis of spectra in the original data cube integrated into different apertures (Sect. 5.1), suggests that these younger stars belong to the galaxy disk. Our SFH of the center of the host galaxy suggests that the latter had later and relatively recent episodes of star formation. This did not happen in the NSC, whose stars were all formed at the early stages of the evolution of M 74. Its SFH shows that its stars were all formed more than 8 Gyr ago. Regarding the origin of these old and metal-poor stars, we cannot say if they were formed in situ a long time ago, when M 74’s gas was still metal-poor. Old ages and low metallicities (although typically lower than what we find here) in NSCs are traditionally explained by globular cluster inspiral (Neumayer2020; Fahrion2021; Fahrion2022a; Fahrion2022b). While globular cluster migration is thought to be the dominant formation mechanism for NSCs less massive than the one in M 74 and hosted by low-mass galaxies, our results do not provide enough information to rule out (or support) this scenario versus a very early in-situ formation (see also Sect. 6.3). Furthermore, the presence of LMXBs in the NSC suggested by the nuclear X-ray emission (see Sect. 3.1), consistent with the NSC old age, provides additional support for the NSC to be one or more globular clusters that migrated to the galactic center.
@@ 6.2 Is M 74’s nuclear star cluster unique?
The NSC in M 74 shows common stellar population properties with other NSCs in the literature. However, it is important to discuss NSC properties in the context of their host galaxy. Fahrion2021 analyzed NSCs in a sample of 25 ETGs in the Fornax and Virgo clusters, finding several of them as old as M 74’s NSC and very metal-poor. However, while the sample includes galaxies of similar stellar masses to M 74 (Sect. 3), the NSCs with the most similar stellar populations to M 74’s NSC were lower-mass NSCs ($\lesssim 10^{6}$ M) hosted by less massive, mostly quiescent ETGs ($\lesssim 10^{9}$ M). ETGs of similar masses as M 74 host metal-rich NSCs. Among the seven dwarf ETGs in the Fornax cluster, analyzed by Johnston2020, NSCs with similar ages to the one in M 74 are much more metal-poor. These NSCs and their dwarf host galaxies are systematically more metal-poor than in the M 74 case. M 74’s NSC age is also similar to three NSCs from Fahrion2022b, who analyzed a sample of nine star-forming dwarf LTGs. While these galaxies are star-forming and have morphological types more similar to M 74, they have much lower stellar masses (${\sim}10^{7}$ to ${\sim}10^{9}$ M). These host galaxies and their NSCs also show systematically lower metallicities than M 74 and its NSC. Some NSCs hosted by galaxies more similar to M 74 in terms of morphological types and masses were analyzed by Kacharov2018. However, while some have similar metallicities to M 74’s NSC (see also Walcher2006), they are all relatively young. While there are other spiral galaxies with NSC ages similar to M 74’s, these galaxies are slightly earlier types than M 74 (Rossa2006).
@@
Figure 7: [Fe/H] metallicity of NSCs as a function of the galaxy stellar mass, reconstructed and updated from Fig. 9 in Neumayer2020. ETGs are indicated in red and LTGs in blue. Red circles are points from spectroscopic and photometric data (Koleva2009; Paudel2011; Spengler2017; Kacharov2018), as from Neumayer2020. The two black solid lines are the mass-metallicity relations of galaxies, given as reference, at the low-mass end (Kirby2013) and the high-mass end (Gallazzi2005). We have added the red triangles, which are NSC total metallicities ([M/H]) from Fahrion2021, and are therefore considered as upper limits for [Fe/H]. We have also added all the blue points. Blue triangles are [M/H], and thus upper limits, of NSCs in dwarf LTGs in Fahrion2022b. Blue circles are NSC [Fe/H] in more massive LTGs in Kacharov2018. Finally, the blue open circle indicates the [Fe/H] metallicity of the decomposed M 74’s NSC (Sect. 4.1.2). For comparison, [Fe/H] of the decomposed host galaxy is indicated with an open square, while a star indicates the [Fe/H] from the original data cube (all extracted from a PSF-sized region, Sect. 5.1).

### v2

@@ 5.1 Stellar populations of the galactic center: the NSC compared with the host galaxy
5.1.1 Average stellar population properties within a PSF aperture
[…]
On the other hand, the difference between the mass-weighted ages of the NSC and its host galaxy is smaller (than when considering light-weighted ages), with the stars in the host galaxy being only about 2 Gyr younger than the NSC. The smaller age difference between the host galaxy and the NSC, in the mass-weighted averages, stems from an older host-galaxy age. This suggests that, while the disk light is dominated by young stars, it also contains a significant (mass) fraction of old populations. However, this is still a significant difference for mass-weighted results, which are expected to be generally biased towards older ages (e.g., Sánchez-Blázquez et al. 2014b; Sattler et al. 2023, 2025). Mass-weighted and light-weighted metallicities are quite similar to each other, while light-weighted [Mg/Fe] abundances are slightly higher than the mass-weighted ones. Uncertainties indicated in Table 1 were obtained via Monte Carlo simulations (Sect. 4.2). Differences in the stellar-population properties of the different components are larger than these uncertainties and thus they can be considered significant. However, these represent only statistical uncertainties and they are relatively low for such high S/N spectra. Systematic uncertainties, such as those due to imperfections in the SSP models, are very challenging to estimate and are not taken into account here.
5.1.2 Additional tests and comparisons
Additionally, we analyzed the ”original” data cube (defined in Sect. 4.1.2 as the prepared data cube just before applying C2D), to assess the robustness of our results. We first compare the results above with the stellar populations obtained from a central PSF aperture in the original data cube (spectrum indicated as a blue solid line in Fig. 4). This contains the integrated light of the NSC and the host galaxy together and, as expected, gives intermediate stellar populations between the NSC and the host galaxy (Table 1). We can also compare the stellar populations in a central PSF aperture with the surrounding region of the galaxy, both extracted from the original data cube. For that, we calculated the average ages, [M/H] and [Mg/Fe] within a much larger central aperture of 100 pc (excluding the central $3\sigma$ PSF to make sure the contamination from the NSC is negligible). For this, we used the stellar-population maps extracted from the original integrated data cube (Appendix A). These stars are younger (5.6 Gyr) and slightly more metal-rich (0.1-0.1 dex), than the stars in the central PSF aperture extracted from the same original data cube (Table 1). However, the [Mg/Fe] abundance is very similar (0.1 dex). The fact that the original data cube, including the light of stars of both the NSC and the host galaxy, shows gradients with older and more metal-poor stars in the very center, is in perfect agreement with the presence of an older and more metal-poor component in the center (i.e. the NSC), which is not present in the surroundings.
We have also tested the robustness of this gradient by fitting a spectrum from the original data cube, but integrated now in a $3\sigma$ PSF (about 40 pc, with a much stronger contribution from the host galaxy). In a $3\sigma$ PSF aperture, we find ages about 1 Gyr younger than integrating within $3\sigma$ PSF, in the original and host-galaxy data cubes, while the NSC age remains the same. Since we are comparing results from slightly different methods (on the one hand, fitting spectra integrated in specific apertures, on the other hand, averaging stellar-population properties in the spatially resolved maps shown in Appendix A), we have checked that the same results were found within $1\sigma$ PSF with the different methods. The mean age, [M/H] and [Mg/Fe] from a $1\sigma$ PSF aperture in the spatially resolved maps in Fig. 8 are respectively: $6.8$ Gyr, 0.2-0.2 and $0.1$ dex, approximately the same as the results from fitting a spectrum from the original data cube obtained integrating in a $1\sigma$ PSF aperture (shown in Table 1).
To additionally test how robust our C2D method is, we have also compared our results with the ones from a NSC spectrum obtained by subtracting the underlying host-galaxy light from a PSF integrated spectrum (”original”, in Fig. 4). Following Fahrion et al. (2021), the host-galaxy spectrum was obtained from an annulus surrounding the NSC, with inner and outer radii of 0.8 ($\sim 3\sigma$ PSF) and 2 arcsec. Before subtracting the host-galaxy spectrum from the ”original” spectrum, the annulus spectrum needs to be first rescaled to the light of the host galaxy in the same central PSF aperture where the ”original” spectrum was integrated (brighter than the annulus). While Fahrion et al. (2021) performed a dedicated morphological decomposition to calculate the scaling factor, we used our decomposition from C2D. We obtained again very old light-weighted ages ($8.5\pm 1.1$ Gyr), low metallicities ($0.39\pm 0.04$ dex) and close to solar [Mg/Fe] enhancement ($0.01\pm 0.03$ dex). While these results give small differences (taking into account the uncertainties - slightly younger ages and higher metallicities), they are closer to the results from C2D than what was obtained from the original data cube. While it is not surprising that different methods give slightly different results (see also Sect. 6.1), our method is more sophisticated than simply subtracting spectra. Our method leads to a cleaner NSC spectrum than the simple subtraction, and therefore to larger differences between results from integrated and decomposed spectra. The main reason for this is that C2D performs a fitting spaxel by spaxel and spectral pixel by spectral pixel, while when subtracting the annulus spectrum from the PSF spectrum we are forced to take an ”average” spectrum within the annulus, and the scaling factor is also calculated using integrated light. More importantly, these differences are not critical for the interpretation of our results: the NSC appears to be very old and metal poor, and much older and more metal poor than the host galaxy according to the results obtained with all methods.
5.1.3 Star formation histories (SFHs) of the NSC and comparison with the host galaxy
We analyze now the mass fraction assigned to SSP models of different ages during the fitting process, which were used to calculate the mass-weighted averages in Table 1. This allows us to reconstruct SFHs in terms of the mass in stars of different age bins. We extracted SFHs of the NSC and the host galaxy from pPXF fits of the spectra extracted in our $1\sigma$ PSF aperture (Sect. 4.2). We show these SFHs in Fig. 5, respectively in the top and middle panels. In the bottom panel we show, as a reference, the SFH of the full non-decomposed galaxy, extracted from the same PSF aperture in the original MUSE data cube before the C2D decomposition (blue spectrum in Fig 4).
The NSC shows distinct stellar populations at old ages, with the highest mass fractions distributed in ages between 10 and 11 Gyr. The host galaxy shows an old component made of two peaks of different ages (${\sim}10$ and ${\sim}12$ Gyr) and a relatively recent component with ages of about 2 Gyr and younger. Some intermediate-age populations are also present (${\sim}6$ Gyr). The SFH of the full galaxy (bottom panel) shows a combination of all these components. A word of caution is deserved by the large uncertainties, indicated by the shades in each panel of Fig. 5. These shades cover the region between the 16% and 84% percentiles, calculated on the distribution of the 100 Monte Carlo realizations. Since mass fractions are normalized to the total mass in each SFH (in each fitted spectrum), these errors indicate that some mass fraction assigned to the highest peaks could be, in reality, redistributed to minor peaks. Overall, the SFHs suggest that the center of the galaxy had an intense early phase of star formation, between 14 and 8 Gyr ago, when the NSC stars formed. Afterwards, it experienced a long quiescent phase, with a very low star-formation rate. Only relatively recently (later than 3 Gyr ago), a more active star formation was reactivated in M 74. However, while this episode involved an extended region of the disk and was enhanced in spiral arms, as shown in Fig. 6, it did not affect the NSC, as this remained old and metal poor. Sánchez-Blázquez et al. (2014a) found a very similar SFH for M 74, using the STEllar Content via Maximum A Posteriori likelihood (STECKMAP) spectral-fitting code (Ocvirk et al. 2006).
5.2 Spatially resolved stellar populations of the host galaxy
Spatially resolved properties of the host galaxy are important to interpret the NSC properties in the context of its surroundings. We mapped the mean age, [M/H] and [Mg/Fe] of the host galaxy, fitting the host-galaxy data cube decomposed with C2D (Sect. 4.1.2 and 4.2). The maps are shown in Fig. 6. Both [M/H] and [Mg/Fe] are presented in logarithmic scale and with respect to solar values (zero values correspond to solar values). The age map (top panel) shows a relatively young host galaxy (with most regions younger than ${\sim}4$ Gyr). A pattern following the spiral arms can be identified, with alternative regions dominated either by young (about ${\sim}2$ Gyr-old or younger) or old stars (older than ${\sim}8$ Gyr). The [M/H] map (middle panel) displays a clear negative radial gradient, with a central metal-rich region of about 500 pc size (supersolar metallicities). Outer regions have alternatively higher and lower values. [Mg/Fe] abundances are mostly anti-correlated with [M/H], showing the lowest values (close to solar) in the central metal-rich ${\sim}500$ pc. Hints of the spiral-arm shape are visible in both [M/H] and [Mg/Fe] maps, whose values do not necessarily follow any correlation or anti-correlation with age. However, the youngest (outer) regions in the analyzed data cube show a strong $\alpha$ enhancement in some metal-poor blobs.
@@ 5.2 Spatially resolved stellar populations of the host galaxy
However, the NSC is as metal-poor as some young regions of the disk. Regarding [Mg/Fe], the NSC has a similar abundance to its surrounding galaxy. These results suggest that while M 74 kept forming stars in most of its disk, this did not happen in some specific regions. While more recent star formation occurred in the central region of the galaxy, allowing for an extended chemical enrichment, this did not happen in the NSC, which remained old and metal-poor.
@@
In Appendix A, we present for reference the analysis of the original MUSE data cube before the spectro-photometric decomposition. Although the Voronoi binning is slightly different, Fig. 8 shows similar trends to Fig. 6. The difference between the original data cube and the host-galaxy decomposed data cube is that the former includes the light of the NSC, which was extracted from the latter. Thus, we can expect a difference in the stellar populations only in the region where the NSC dominates. Indeed, as described in Appendix A, a clear [M/H] drop is visible within a PSF size in Fig. 8. Our results for the central region of M 74 are in agreement with the stellar-population analysis by Pessa et al. (2023, see also Sect. 3).
@@ 6.1 The nuclear star cluster in M 74 and its formation
Our stellar population analysis has shown an ancient and metal-poor NSC, surrounded by a much younger and more metal-rich region of the host galaxy. As shown by JWST observations (Hoyer et al. 2023b), the NSC lives in a cavity with no gas or dust, while being hosted by an actively star-forming galaxy with a complex structure of gas and dust filaments. The cavity formed due to some processes either keeping gas and dust out of it or wiping them out of the central region (see also Sect. 6.3). While the presence of a massive black hole in M 74 is still under debate, the presence of a low-luminosity AGN was proposed by different authors (e.g., Dong and De Robertis 2006; She et al. 2017). It is also one of the possible explanations for the presence of the cavity (all gas may have been blown via AGN feedback, especially if the black-hole activity was much stronger in the past) and of the mid-infrared emission found at the center of the galaxy with JWST observations (Hoyer et al. 2023b). Similarly, central quiescent regions (of about 100 pc sizes), surrounded by starbursts, as well as nuclear older ages (within radii from few to ${\sim}50$ pc) than the circumnuclear regions, were previously detected in AGN hosts (e.g., Márquez et al. 2025; da Silva et al. 2025).
In any case, the presence of the cavity may have affected the evolution of the NSC, which could not grow via in-situ star formation due to the lack of gas. However, the surrounding region of the NSC, including stars in the region of the cavity, is much younger and more metal-rich. Our decomposition, supported by the analysis of spectra in the original data cube integrated into different apertures (Sect. 5.1.2), suggests that these younger stars belong to the galaxy disk. Our SFH of the center of the host galaxy suggests that the latter had later and relatively recent episodes of star formation. This did not happen in the NSC, whose stars were all formed at the early stages of the evolution of M 74. Its SFH shows that its stars were all formed more than 8 Gyr ago. Regarding the origin of these old and metal-poor stars, we cannot say if they were formed in situ a long time ago, when M 74’s gas was still metal-poor. Old ages and low metallicities (although typically lower than what we find here) in NSCs are traditionally explained by globular cluster inspiral (Neumayer et al. 2020; Fahrion et al. 2021, 2022b, 2022a). While globular cluster migration is thought to be the dominant formation mechanism for NSCs less massive than the one in M 74 and hosted by low-mass galaxies, our results do not provide enough information to rule out (or support) this scenario versus a very early in-situ formation (see also Sect. 6.3). Furthermore, the presence of LMXBs in the NSC suggested by the nuclear X-ray emission (see Sect. 3.1), consistent with the NSC old age, provides additional support for the NSC to be one or more globular clusters that migrated to the galactic center.
@@ 6.2 Is M 74’s nuclear star cluster unique?
The NSC in M 74 shows common stellar population properties with other NSCs in the literature. However, it is important to discuss NSC properties in the context of their host galaxies. Fahrion et al. (2021) analyzed NSCs in a sample of 25 ETGs in the Fornax and Virgo clusters, finding several of them as old as M 74’s NSC and very metal-poor. However, while the sample includes galaxies of similar stellar masses to M 74 (Sect. 3), the NSCs with the most similar stellar populations to M 74’s NSC were lower-mass NSCs ($\lesssim 10^{6}$ M) hosted by less massive, mostly quiescent ETGs ($\lesssim 10^{9}$ M). ETGs of similar masses as M 74 host metal-rich NSCs. Among the seven dwarf ETGs in the Fornax cluster, analyzed by Johnston et al. (2020), NSCs with similar ages to the one in M 74 are much more metal-poor. These NSCs and their dwarf host galaxies are systematically more metal-poor than in the M 74 case. M 74’s NSC age is also similar to three NSCs from Fahrion et al. (2022a), who analyzed a sample of nine star-forming dwarf LTGs. While these galaxies are star-forming and have morphological types more similar to M 74, they have much lower stellar masses (${\sim}10^{7}$ to ${\sim}10^{9}$ M). These host galaxies and their NSCs also show systematically lower metallicities than M 74 and its NSC. Some NSCs hosted by galaxies more similar to M 74 in terms of morphological types and masses were analyzed by Kacharov et al. (2018). However, while some have similar metallicities to M 74’s NSC (see also Walcher et al. 2006), they are all relatively young. While there are other spiral galaxies with NSC ages similar to M 74’s, these galaxies are slightly earlier types than M 74 (Rossa et al. 2006).
@@
Figure 7: NSC metallicity as a function of the galaxy stellar mass (figure reconstructed and updated from Fig. 9 in Neumayer et al. (2020)). The blue open circle indicates the [Fe/H] metallicity of the decomposed M 74’s NSC. For comparison, [Fe/H] of the decomposed host galaxy is indicated with an open square, while a star indicates the [Fe/H] from the original (integrated) data cube of M 74 (see the text for all the details). ETGs are indicated in red and LTGs in blue. Red circles are points from spectroscopic and photometric data (Koleva et al. 2009; Paudel et al. 2011; Spengler et al. 2017; Kacharov et al. 2018), as from Neumayer et al. (2020). The two black solid lines are the mass-metallicity relations of galaxies, given as reference, at the low-mass end (Kirby et al. 2013) and the high-mass end (Gallazzi et al. 2005). We have added the red triangles, which are NSC total metallicities ([M/H]) from Fahrion et al. (2021), and are therefore considered as upper limits for [Fe/H]. We have also added all the blue points. Blue triangles are [M/H], and thus upper limits, of NSCs in dwarf LTGs in Fahrion et al. (2022a). Blue circles are NSC [Fe/H] in more massive LTGs in Kacharov et al. (2018).